Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.
i'm not sure how to start and cant find any problem simlar to this.

Found 2 solutions by Alan3354, lwsshak3:
Answer by Alan3354(69443)

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.
-------------
y = –x + 6 is in slope-intercept form, y = mx + b, where m = slope and b = the y-intercept.
----------
The slope of y = –x + 6 is -1
Parallel lines have the same slope, so the line thru (-4,2) will have a slope of -1 also.
Use y = mx + b and the point to find b
2 = -1*(-4) + b
b = -2
-------
--> y = -x - 2

Answer by lwsshak3(11628)

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form
**
Standard form of equation for straight line: y=mx+b, m=slope, b=y-intercept
y=–x+6
The slope, m, of this given equation=-1
An equation parallel to given equation would have the same slope, but the y-intercept may be different.
y=-x+b
solving for b using (x,y) coordinates from given point (-4,2)
2=-(-4)+b
2=4+b
b=-2
equation:
y=-x-2s